OpenStudy (anonymous):

solve the system by using elimination or elimination with multiplication: 3x - 2y = 9 and -3x + 2y = -9.

OpenStudy (anonymous):

3x - 2y = 9 -3x + 2 = -9 is there a y next to the 2 on the second equation?

OpenStudy (anonymous):

yes, sorry it should be 3x - 2y = 9 and -3x + 2y = -9. thank you

OpenStudy (anonymous):

if it is like this: 3x - 2y = 9 -3x + 2y = -9 then In order to solve for one variable, we must eliminate the other variable. So if we wanted to solve for y, we would have to eliminate x (or vice versa). So lets eliminate x. In order to do that, we need to have both x coefficients that are equal but have opposite signs (for instance 2 and -2 are equal but have opposite signs). This way they will add to zero. So to make the x coefficients equal but opposite, we need to multiply both x coefficients by some number to get them to an equal number. So if we wanted to get 3 and -3 to some equal number, we could try to get them to the LCM. Since the LCM of 3 and -3 is -3, we need to multiply both sides of the top equation by -1 and multiply both sides of the bottom equation by -1 like this: -1(3x - 2y) = 9 * -1 Multiply the top equation (both sides) by -1 -1(-3x + 2y) = -9 * -1 Multiply the bottom equation (both sides) by -1 So after multiplying we get this: -3x + 2y = -9 3x - 2y = 9 Notice how -3 and 3 add to zero, 2 and -2 add to zero, -9 and 9 and to zero. So in essence, they cancel each other out. So we're left with 0=0 which means any x or y value will satisfy the system of equations. So there are an infinite number of solutions. So this system is dependent.

OpenStudy (anonymous):

Do you get this?

OpenStudy (anonymous):

yes, thank you so much.